OpenStudy (becca9898):

What is an example of an equation with 2 real zeros?

OpenStudy (kira_yamato):

x(x - 1) = 0

hartnn (hartnn):

becca, can you make such equation on your own now? ask doubts, you'll learn more...

OpenStudy (becca9898):

I have tried to come up with an equation. I'm homeschooled and my lessons do not add much detail to help me learn.

hartnn (hartnn):

oh, you tried, good what did u have ?

OpenStudy (becca9898):

I have tried the equation -4x^2+3x because the equation also has to be an upward opening parabola

hartnn (hartnn):

in \(ax^2+bx+c=0\) if a is positive, then parabola opens upward if a is negative then parabola opens downwards. NOTE : -4x^2+3x is an EXPRESSION and EQUATION has "=" sign, like -4x^2+3x =0 or -4x^2+3x =-10

hartnn (hartnn):

so, try again to get equation with upward opening parabola ?

OpenStudy (becca9898):

The leading coefficient then has to become positive, so 4x^2+3x works

hartnn (hartnn):

but thats an expression where is the equation ?

OpenStudy (becca9898):

That's where I am confused. It's a story problem that is in three parts and none of it makes any sense

OpenStudy (becca9898):

I'm not sure what real zeros are because they didn't mention them until now.

hartnn (hartnn):

you know what real numbers are ?

hartnn (hartnn):

and i was just asking for equation, 4x^2+3x =0 will work as an equation with parabola openeing upwards and real zeros

OpenStudy (becca9898):

How do I know there are 2 real zeros?

hartnn (hartnn):

a polynomial with degree 'n' has 'n' zeros. does this make sense ? know what is a degree of polynomial ?

hartnn (hartnn):

degree is the highest exponent of 'x' you knew this ?

OpenStudy (becca9898):

The degree would be a 3rd degree trinominal, correct?

OpenStudy (becca9898):

Oops no. 3rd degree binomial

hartnn (hartnn):

how do you find degree ?

hartnn (hartnn):

what is the highest exponent of 'x' in \(ax^2+bx+c=0\) ?

OpenStudy (becca9898):

I learned that you add the exponets together for the degree... I'm assuming that is wrong.

OpenStudy (becca9898):

Would it be 2?

hartnn (hartnn):

yes! and if there are more than 1 variables, then you ADD exponents, but here there's only 1 variable , 'x' so, degree = 2, so it MUST HAVE 2 'zeros' like in \(ax^7+bx^2+x+1=0\), the highest exponent = degree = 7 so this will have 7 'zeros' understanding this ?

OpenStudy (becca9898):

This is making a lot more sense!!!

OpenStudy (becca9898):

When it says an upward parabola does it mean the vertex is pointing up or down?

hartnn (hartnn):

nice!! so, when you were asked to have '2' zeros, you will make an equation with highest exponent = 2 so, use ax^2+bx+c = 0 , put values of a,b,c but ax^4+bx^2+d =0 will not work because degree =4 and it will have more than 2 zeros

hartnn (hartnn):